# Emulating Full Client Participation: A Long-Term Client Selection Strategy for Federated Learning

### v1

@@ V-C Individual Fairness Analysis
We analyze the proposed individual fairness in terms of the trade-off factor $V$, the similarity measure $\epsilon$, and the probability difference measure $\delta$. The results are illustrated using the 2SPC scenario on the CIFAR-10 dataset, and similar trends are observed for other scenarios. The random selection strategy is chosen as the baseline because it selects clients randomly and without bias. The impact of these parameters is assessed through two metrics: test accuracy and the standard deviation of the selection probability $p_{i}$ between clients with similar data distributions, denoted as $\sigma(t)$. To calculate $\sigma(t)$, we define $c_{i}(t)$ as the accumulated number of selections for client $i$ during previous $t$ communication rounds. That is, $c_{i}(t)=\sum_{k=1}^{t}x_{i,k}$. For each client $i$, we identify clients whose similarity with it is less than $\epsilon$, denoted as $\mathbb{I}_{i}(t)=\{j\in\mathbb{N}\mid\text{Dist}_{i,j}(t)<\epsilon\}$. We then compute the average selection count among clients in $\mathbb{I}_{i}(t)$ by
@@
Trade-Off Factor $V$. We vary the value of $V$ from $0.0$, $0.4$, $0.6$, and $1.0$, and present the results in Fig. 6. First, we observe that the proposed method with these four values converges faster than the random strategy and achieves better performance. Among them, $V=0.6$ yields the best test accuracy. This suggests that the individual fairness constraint improves the overall performance. Second, the random strategy achieves the lowest standard deviation $\sigma(t)$. The proposed method exhibits significantly larger $\sigma(t)$ when $V=1.0$. This is because when $V=1.0$, our proposed method prioritizes minimizing the estimation error in Eq. (2) and neglects the fairness constraint. Furthermore, the standard deviation $\sigma(t)$ decreases significantly when $V\leq 0.6$. It indicates that the constraint in Eq. (7) indeed enhances the system fairness.
Similarity Measure $\epsilon$. We vary the value of $\epsilon$ across $0.05$, $0.1$, $0.2$, and $0.3$, and observe the corresponding test accuracy and standard deviation, as presented in Fig. 7. First, the results on test accuracy indicate that the proposed method outperforms the random strategy consistently across these $\epsilon$ values. On the other hand, a larger $\epsilon$ in the proposed method corresponds to a smaller standard deviation. Specifically, the results are comparable to the random selection strategy when $\epsilon>0.05$, and even surpass it when $\epsilon=0.3$. This is attributed to the fact that a larger $\epsilon$ enforces more clients to follow the fairness constraint, thus the standard deviation obtained through the proposed strategy gets closer to the random selection strategy.
@@
Probability Difference Measure $\delta$. We vary the value of $\delta$ across $0.01$, $0.1$, $0.15$, and $0.2$, and observe the corresponding test accuracy and standard deviation, as presented in Fig. 8. The results indicate that the proposed method consistently performs better than the random selection strategy across these $\delta$ values in terms of test accuracy. Moreover, a smaller $\delta$ in the proposed method corresponds to a smaller standard deviation. This is because a smaller $\delta$ narrows the range of probability difference between clients with similar distributions, resulting in a smaller standard deviation.
@@ VI Conclusion
In our work, we introduce a novel long-term client selection approach for federated learning, aiming to emulate the optimization trajectory achieved by full client participation. We establish two guiding principles to steer the client selection process effectively. We formulate the client selection problem as a long-term optimization task, where the objective is to minimize the estimation error in gradient space between the selected subset of clients and the full client set. Besides, we impose the individual fairness constraint to ensure similar selection frequencies among clients with similar data distributions. Leveraging Lyapunov optimization techniques from control theory, we transform this problem and propose a greedy algorithm to find solutions efficiently. Experimental results demonstrate that our proposed strategy closely approximates the performance achieved under full client participation, with significant improvements over previous methods in terms of both accuracy and fairness. Additionally, our method exhibits efficiency by incurring minimal time overhead, which is often negligible when compared to local training duration.
Figure 8: Impact of the probability difference measure $\delta$.

### v2

@@ V-B2 Fairness
We evaluate the fairness through the standard deviation in the selection probability, denoted as $\sigma(t)$. To calculate $\sigma(t)$, we define $c_{i}(t)$ as the accumulated number of selections for client $i$ during previous $t$ rounds. That is, $c_{i}(t)=\sum_{k=1}^{t}x_{i,k}$. For each client $i$, we identify clients whose similarity with it is less than $\epsilon$, denoted as $\mathbb{I}_{i}(t)=\{j\in\mathbb{N}\mid\text{Dist}_{i,j}(t)<\epsilon\}$. We then compute the average selection count among clients in $\mathbb{I}_{i}(t)$ by
@@
We compare the proposed method with three baselines: AFL [10], Powerd [11], and divFL [16], along with their respective versions incorporating the uniform selection constraint, denoted as AFL+Fair, Powerd+Fair, and divFL+Fair. These methods are evaluated based on two key aspects: test accuracy, and the standard deviation $\sigma(t)$. The analysis is conducted on both the CIFAR-10 and FMNIST datasets under the 2SPC scenario, with similar trends observed in other settings. The results of the CIFAR-10 and FMNIST datasets are illustrated in Fig. 4 and Fig. 5, respectively. In these figures, the vanilla methods are represented by solid lines, while the methods incorporating the uniform selection constraint are indicated by dashed lines.
TABLE I: Time overhead analysis on FMNIST in the 1SPC scenario.
Method Random AFL Powerd ShapleyFL FedRank divFL LongFed
Time (ms) for client selection 0.038 0.153 0.016 221.023 0.057 0.219 0.529
Time (ms) for a complete round 2019 2044 2033 3242 2045 2407 2472
In Fig. 4, the test accuracy results indicate that the proposed method achieves the best performance. However, applying the uniform selection constraint leads to a degradation in model performance for all three baselines. On the other hand, from the perspective of fairness, incorporating the uniform selection constraint reduces the standard deviation $\sigma(t)$ for all three baseline methods. Notably, divFL+Fair achieves the lowest $\sigma(t)$ among all methods. The proposed methods attains a comparable but slightly higher standard deviation than divFL+Fair. This phenomenon occurs because the uniform selection constraint enforces equal selection probabilities across all clients, whereas the proposed method ensures similar selection probabilities only for clients with similar data distributions. As a result, divFL+Fair exhibits slightly better fairness performance than our method. Considering both test accuracy and fairness (as measured by $\sigma(t)$), the proposed individual fairness approach demonstrates a superior balance, effectively improving both model performance and fairness simultaneously compared to the uniform selection constraint.
Figure 5: Fairness results on the FMNIST dataset in the 2SPC scenario.
In Fig. 5, the results for the standard deviation $\sigma(t)$ follow a similar trend to that of the CIFAR-10 dataset. That is, introducing the uniform selection constraint improve fairness, and the proposed LongFed achieves fairness comparable to divFL+Fair. For test accuracy, we observe that applying the uniform selection constraint leads to a decline in performance for Powerd, while AFL maintains similar performance, and divFL exhibits a slight improvement. This may be because these methods typically select the same subset of clients across multiple rounds, potentially leading the system to a suboptimal solution. The introduction of the uniform selection constraint forces these methods to select different clients, particularly those that were previously under-selected, helping to escape the suboptimal state. Most importantly, the proposed LongFed still achieves the best overall performance, demonstating its effectivess in balancing both fairness and model accuracy.
V-B3 Time Overhead Analysis
We evaluate the time overhead of the proposed client selection strategy, and the results are presented in Table I. The analysis is conducted on the FMNIST dataset under the 1SPC scenario, with similar trends observed across other scenarios. In Table I, the first row highlights the time dedicated solely to client selection, while the second row denotes the overall time required for a complete round, including client selection, local updates, and global aggregation.
First, examining the time specifically for client selection, our proposed method exhibits only a marginal increase (less than 0.4ms) compared to existing methods. This slight increase is primarily due to the computational cost of evaluating the distance $\text{Dist}_{i,j}(t)$ in Eq. (3). However, this can be mitigated by employing more efficient distance computation techniques in high-dimensional gradient space. More importantly, the time required for client selection (approximately 0.5ms) is negligible compared to the total time required to complete a round (approximately 2000ms). Therefore, the marginal increase in time for the proposed strategy is justified, given its superior improvements in model performance and fairness.
Additionally, ShapleyFL [22] incurs significantly higher selection time due to the computationally expensive process of Shapley values, which involves combinatorial computations.
V-C Visualization of Client Selection Strategy
Figure 6: Visualization of the selected clients on the FMNIST dataset under the 1SPC scenario.
Figure 7: Visualization of the selected clients on the FMNIST dataset under the 2SPC scenario.
We provide visualizations of selection results to offer an interpretable analysis. In Fig. 6, (a) is the t-SNE plot of client embeddings, where clients are organized into 10 clusters. (b)-(j) display the results of the selected clients using different client selection strategies, with the chosen clients marked by black stars. As shown in Fig. 6 (b)-(f), baseline methods often select two or more clients from the same cluster. In contrast, as shown in Fig. 6 (i), the proposed LongFed selects one client from each cluster, effectively approximating the data distribution of the full client set. Additionally, as shown in Fig. 6 (g) and (h), divFL tends to select the same set of clients across successive rounds. In contrast, as shown in Fig. 6 (i) and (j), the proposed LongFed selects a more diverse subset of clients across multiple rounds, ensuring fairness in multi-round selection.
Results in the 2SPC scenario are illustrated in Fig. 7. (a) shows the t-SNE plot of client embeddings, where the clustering pattern is not evident. In (b)-(j), circles are drawn around the selected clients, indicating that clients covered by a circle can be represented by the corresponding selected client. For baseline methods, as shown in Fig. 7 (b)-(e), the circles around selected clients often overlap and cover only a minority of clients. In our proposed method, as shown in Fig. 7 (i), these circles cover the majority of clients. This suggests that the selected clients provide a better approximation of the full client set compared to the baseline approaches.
both FedRank and divFL repeatedly select the same subset of clients. Notably, this subset remains unchanged after 55 rounds and continues to be selected across nearly 250 rounds. As a result, the global model is trained on a limited set of clients, leading to biased predictions for the remaining clients. This issue is corroborated by the test accuracy results in Fig.3 (b), where divFL and FedRank exhibit the worst performance. Additionally, as shown in Fig. 5, divFL has the highest standard deviation $\sigma(t)$ among the baselines, further highlighting its fairness limitations. In contrast, LongFed maintains a more diverse client selection over multiple rounds, ensuring both fairness and improved generalization. These findings are consistent with the observations in the 1SPC scenario. Additional visualizations for other scenarios are provided in the supplementary file.
Moreover, as shown in Fig. 7 (f)-(h), FedRank and divFL select the same subset of clients. Notably, this subset remains unchanged after 55 rounds and continues to be selected across nearly 250 rounds for both FedRank and divFL. As a result, the global model is trained only on the 5 selected clients, leading to biased predictions for the remaining clients. This issue aligns with the test accuracy results in Fig.3 (b), where divFL and FedRank exhibit the worst performance. Additionally, as shown in Fig. 5, divFL has the highest standard deviation $\sigma(t)$ among the baselines, further highlighting its fairness limitations. In contrast, LongFed maintains a more diverse client selection over multiple rounds, ensuring both fairness and improved generalization. These findings are consistent with the observations in the 1SPC scenario. Additional visualizations for other scenarios are provided in the supplementary file.
V-D Individual Fairness Analysis
We analyze the $\epsilon\text{-}\delta$-individual fairness in terms of the trade-off factor $V$, the similarity measure $\epsilon$, and the probability difference measure $\delta$. The results are illustrated using the 2SPC scenario on the CIFAR-10 dataset, and similar trends are observed for other scenarios. We consider two metrics: test accuracy, and the standard deviation $\sigma(t)$ in Eq. (34).
@@ V-D Individual Fairness Analysis
Trade-Off Factor $V$. We vary the value of $V$ from $0.0$ to $1.0$, where a smaller $V$ indicates a higher priority for the individual fairness constraint. The results are presented in Fig. 8. First, we observe that the proposed method achieves the highest test accuracy when $V=0.6$ and $V=0.8$. Second, in terms of fairness, the proposed method exhibits significantly larger standard deviation $\sigma(t)$ when $V=1.0$. This is because when $V=1.0$, the method solely prioritizes minimizing the estimation error in Eq. (2) while neglecting the fairness constraint. Furthermore, the standard deviation $\sigma(t)$ decreases significantly when $V\leq 0.8$. Considering both model performance and fairness, we recommend setting $V=0.8$ as the optimal choice.
@@
Similarity Measure $\epsilon$. We vary the value of $\epsilon$ across $0.05$, $0.1$, $0.2$, and $0.3$, and evaluate the corresponding test accuracy and standard deviation, as shown in Fig. 9. First, we observe that test accuracy remains relatively consistent and is not particularly sensitive to the choice of $\epsilon$. However, a larger $\epsilon$ results in a smaller standard deviation. This is because $\epsilon$ defines the similarity threshold for grouping clients, and a larger value considers more clients as similar, leading to reduced variability in selection. Considering both model performance and fairness, we recommend setting $\epsilon=0.3$ as the optimal choice.
Figure 10: Impact of the probability difference measure $\delta$.
Probability Difference Measure $\delta$. We vary the value of $\delta$ across $0.01$, $0.1$, $0.15$, and $0.2$, and evaluate the corresponding test accuracy and standard deviation, as shown in Fig. 10. The results indicate that test accuracy remains relatively stable and is not sensitive to the choice of $\delta$. However, a smaller $\delta$ leads to a lower standard deviation. This is because $\delta$ defines the allowable difference in selection probabilities between clients with similar distributions. A smaller $\delta$ enforces a stricter constraint, resulting in a smaller standard deviation. Given the balance between model performance and fairness, we recommend setting $\delta=0.01$ as the best choice.
V-E Summary
In summary, the proposed LongFed demonstrates faster convergence and superior test accuracy, effectively enhancing both model performance and fairness simultaneously. Moreover, LongFed consistently selects representative clients to approximate full participation across multiple rounds, regardless of whether a clustering pattern exists among clients.
@@ VI Conclusion
In this work, we focus on the client selection problem in federated learning, and propose an effective and fair selection method to improve both model performance and fairness. To achieve this, we introduce two guiding principles and formulate the client selection problem as a long-term optimization task. Experiments show that our method effectively guides the system to converge along a trajectory similar to that of full client participation. Visualization results further illustrate that our approach increases data diversity by selecting clients based on their data distributions, thereby improving both model performance and fairness.